Also Present:
D Norman, C Kaleida, A Saha, R Blum
Problems:
4MAP back up at start of the night. Not clear to us what was happening last night, but seems system came up fine today.
At start of night, TCS application froze and had to be restarted. TCS server was ok.
See strange ramp on some images. See DECam_00149355.fits.fz. Images before and after look ok. See trails on many of these images (3003). As if something is out of phase with scripts (i.e. exposures vs telescope moves).
Lose several images in 3002, 9999 because of interlock failures: OCS_STATUS "state=red". We reset to recover, but image in sequence is lost.